OpenStudy (anonymous):

Which polynomial represents the ordered pairs ' (1, 6); (2, 11); (4, 39); and (7, -54).

OpenStudy (anonymous):

i got x3 - 4x2 + 10x - 1 = f(x)

OpenStudy (anonymous):

if you plug in those ordered pairs, do they all work? if they do, that polynomial represents the ordered pairs

OpenStudy (anonymous):

how why would i plug in the ordered pairs

OpenStudy (anonymous):

you plug in the first number from the coordinate into the x's, solve the polynomial (the side thats not f(x)), and if you get the second coordinate as the answer, the ordered pair works example: 2x+3=f(x) ordered pair =(3,5) 2(3)+3 this doesnt equal 5, so the ordered pair doesnt work

OpenStudy (anonymous):

so i got it wrong dang

OpenStudy (anonymous):

yes and its possible theres more than one answer ill try to find one

OpenStudy (anonymous):

was this a multiple choice question?

OpenStudy (anonymous):

x3 + 8x2 + 10x - 1 = f(x) -2x3 - 17x2 - 23x + 32 = f(x) -2x3 + 17x2 - 32x + 23 = f(x) x3 - 4x2 + 10x - 1 = f(x)